# Pushing Cops and Robber on Graphs of Maximum Degree 4

### v1

@@ 3. Graphs with Maximum Degree 4
      Let $U=\left\{v^{\prime},v,u,x^{\prime},x,y^{\prime},y,x_{2},x_{1},y_{2},w_{1},\ldots,w_{8}\right\}$. We note that it is possible that all vertices in $\left\{v^{\prime},v,u,x^{\prime},x,y^{\prime},y,x_{2},x_{1},y_{2},w_{1},\ldots,w_{8}\right\}$ are not distinct. First, we argue that if not vertex of $U$ is pushed and $\mathcal{R}$ moves to either of $x$ or $y$, it will be trapped in at most two cop moves. To see this, let $\mathcal{R}$ moves to $x$ (resp. $y$). Then, $\mathcal{C}$ pushes $x_{2}$ (resp. $y_{2}$), which forces $\mathcal{R}$ to move to $x_{1}$ and since $N^{+}(x_{1})=1$, $\mathcal{R}$ will be trapped in the next cop move by pushing $w_{4}$. Now, as long as $\mathcal{R}$ does not move to $x$ or $y$ (i.e., passes its moves by staying at $x$), $\mathcal{C}$ does the following: Let $\mathcal{C}$ occupies a vertex $z$. $\mathcal{C}$ finds a shortest path in $G$ from $z$ to a vertex in $U$ and either move towards it if the orientation allows or pushes its current vertex so that the orientation in the next cop move allows cop to move towards $U$. Using this strategy, $\mathcal{C}$ will reach a vertex of $U$ without ever pushing any vertex of $U$. During all these moves, if $\mathcal{R}$ moves from $u$, it will be trapped. Hence, $\mathcal{R}$ is still at $u$. Now, consider the possible vertices where $\mathcal{C}$ arrives in $U$. If it arrives at either of $v,v^{\prime}$, then observe that $\mathcal{R}$ will have to move in the next round, else it will be captured and since we have not pushed any vertex in $U$, $\mathcal{R}$ will be trapped in at most two rounds. If $\mathcal{C}$ reaches $x^{\prime}$ (resp., $y^{\prime}$), then $\mathcal{C}$ will push $y$ (resp., $x$), forcing $\mathcal{R}$ to move to $x$ (resp. $y$) in the next round, where it will be captured in the next round. Next, if $\mathcal{C}$ reaches a vertex in $\left\{w_{4},w_{5}\right\}$, then $\mathcal{C}$ pushes $y$ to force $\mathcal{R}$ to move to $x$, and then push $x_{2}$ to force $\mathcal{R}$ to $x_{1}$. If $\mathcal{C}$ is at $w_{5}$, then $\mathcal{R}$ will be captured, else, if $\mathcal{C}$ is at $w_{4}$, then $\mathcal{C}$ pushes $y$, and if $\mathcal{R}$ moves to $w_{4}$ it gets captured, else $\mathcal{C}$ pushes $w_{4}$ to trap $\mathcal{R}$. Finally, if $\mathcal{C}$ reached a vertex in $\left\{w_{1},w_{2},w_{3}\right\}$ (resp. in $\{w_{6},w_{7},w_{8}\}$), then $\mathcal{C}$ pushes $y$ (resp. $x$). This forces $\mathcal{R}$ to move to $x$ (resp. $y$). Next, $\mathcal{C}$ pushes $x_{1}$, forcing $\mathcal{R}$ to move to $x_{2}$ (resp. $y_{2}$). Now, if $\mathcal{C}$ was at $w_{1}$ (resp. $w_{8}$), $\mathcal{R}$ would be captured in this round. Else, without loss of generality, let us assume that $\mathcal{C}$ is at $w_{2}$ (resp. $w_{6}$). In this case, $\mathcal{C}$ pushes $w_{3}$ (resp. $w_{7}$). Now, if $\mathcal{R}$ does not move to $w_{2}$ (resp. $w_{6}$) in the next round, it will be trapped by pushing $w_{2}$ (resp. $w_{6}$), and if it moves, observe that it will be captured by $\mathcal{C}$. This completes our proof for this case.
The proof of the claim is completed by above two exhaustive cases. ∎
The proof of our lemma follows from 3.4 and 3.6. ∎
@@
The cop $\mathcal{C}$ follows the strategy from Lemma 3.3 to ensure the invariant that the out-degree of every visited vertex is at most one. If this invariant breaks, then observe that $\mathcal{C}$ traps $\mathcal{R}$ using Lemma 3.3. Since $\overrightarrow{G}$ is finite, after a finite number of rounds, $\mathcal{R}$ will again visit a visited vertex where it will be trapped using Lemma 3.1. Finally, $\mathcal{C}$ can capture the trapped robber using Proposition 2.1. This completes our proof ∎
Observe that Lemma 1.4 along with Theorem 1.2 and Proposition 1.1, implies the following theorem.
@@ 4. Acknowledgments
This research was supported by the IDEX-ISITE initiative CAP 20-25 (ANR-16-IDEX-0001), the International Research Center “Innovation Transportation and Production Systems” of the I-SITE CAP 20-25, and the ANR project GRALMECO (ANR-21-CE48-0004).

### v2

@@ 4. Graphs with Maximum Degree 4
      Now, as long as $\mathcal{R}$ stays at the vertex $u$ (i.e., it passes its moves by staying at the same vertex), $\mathcal{C}$ does the following: Let $\mathcal{C}$ occupy a vertex $z$. $\mathcal{C}$ finds a shortest path in $G$ from $z$ to a vertex in $U$ and either moves towards it if the orientation allows or pushes its current vertex so that the orientation in the next cop move allows the cop to move towards $U$. Using this strategy, $\mathcal{C}$ will reach a vertex of $U$ without ever pushing any vertex of $U$. During all these moves, if $\mathcal{R}$ moves from $u$, it will be trapped. Hence, $\mathcal{R}$ is still at $u$. Now, consider the possible vertices where $\mathcal{C}$ arrives in $U$. If it arrives at either of $v,v^{\prime}$, then observe that $\mathcal{R}$ will have to move in the next round, else it will be captured and since we have not pushed any vertex in $U$, $\mathcal{R}$ will be trapped in at most two rounds. If $\mathcal{C}$ reaches $x^{\prime}$ (resp. $y^{\prime}$), then $\mathcal{C}$ will push $y$ (resp. $x$), forcing $\mathcal{R}$ to move to $x$ (resp. $y$) in the next round, where it will be captured in the next round. Next, if $\mathcal{C}$ reaches a vertex in $\left\{w_{4},w_{5}\right\}$, then $\mathcal{C}$ pushes $y$ to force $\mathcal{R}$ to move to $x$, and then push $x_{2}$ to force $\mathcal{R}$ to $x_{1}$. If $\mathcal{C}$ is at $w_{5}$, then $\mathcal{R}$ will be captured, else, if $\mathcal{C}$ is at $w_{4}$, then $\mathcal{C}$ pushes $y$, and if $\mathcal{R}$ moves to $w_{4}$ it gets captured, else $\mathcal{C}$ pushes $w_{4}$ to trap $\mathcal{R}$. Finally, if $\mathcal{C}$ reached a vertex in $\left\{w_{1},w_{2},w_{3}\right\}$ (resp. in $\{w_{6},w_{7},w_{8}\}$), then $\mathcal{C}$ pushes $y$ (resp. $x$). This forces $\mathcal{R}$ to move to $x$ (resp. $y$). Next, $\mathcal{C}$ pushes $x_{1}$, forcing $\mathcal{R}$ to move to $x_{2}$ (resp. $y_{2}$). Now, if $\mathcal{C}$ was at $w_{1}$ (resp. $w_{8}$), $\mathcal{R}$ would be captured in this round. Else, without loss of generality, let us assume that $\mathcal{C}$ is at $w_{2}$ (resp. $w_{6}$). In this case, $\mathcal{C}$ pushes $w_{3}$ (resp. $w_{7}$). Now, if $\mathcal{R}$ does not move to $w_{2}$ (resp. $w_{6}$) in the next round, it will be trapped by pushing $w_{2}$ (resp. $w_{6}$), and if it moves, observe that it will be captured by $\mathcal{C}$. This completes our proof for this case.
The proof of the claim is completed by the above two exhaustive cases. ∎
The proof of our lemma follows from ˜4.4 and ˜4.6. ∎
@@
The cop $\mathcal{C}$ follows the strategy from Lemma˜4.3 to ensure the invariant that the out-degree of every visited vertex is at most one. If this invariant breaks, then observe that $\mathcal{C}$ traps $\mathcal{R}$ using Lemma˜4.3. Since $\overrightarrow{G}$ is finite, after a finite number of rounds, $\mathcal{R}$ will again visit a visited vertex where it will be trapped using Lemma˜4.1. Finally, $\mathcal{C}$ can capture the trapped robber using Proposition˜2.1. This completes our proof. ∎
Observe that Lemma˜1.4 along with Corollary˜1.3 and Proposition˜1.2, implies the following theorem.
@@
5. Conclusion
We established that if $\overrightarrow{G}$ is an orientation of a graph with maximum degree $4$, then ${\sf c_{sp}}(\overrightarrow{G})=1$. We also observed that if $\overrightarrow{G}$ can be transformed into a DAG using the push operation, then ${\sf c_{sp}}(\overrightarrow{G})=1$. Huang, MacGillivray, and Yeo [huang2002pushing] showed that a chordal graph can be made a DAG via push operations if and only if it does not contain the subdigraphs depicted in Figure˜5. Consequently, the class of chordal digraphs excluding these subdigraphs is cop-win when the cop has strong push ability. Moreover, Das et al. [das2023cops] proved that if $\overrightarrow{G}$ is an orientation of an interval graph, then ${\sf c_{sp}}(\overrightarrow{G})=1$. This naturally raises the open question of whether all orientations of all chordal graphs are cop-win under strong push ability. We suspect this to be the case.
Figure 5. A chordal digraph can be pushed to be a DAG iff it does not contain any of the two orientations of $K_{4}$ as subdigraphs.
As discussed in Section˜1, triangle-free planar graphs are strong push cop-win. One may wonder if all planar graphs are cop-win when the cop has the strong push ability. Another interesting direction is to extend our result to $4$-degenerate graphs or graphs with maximum degree $5$.
While a broad class of graphs is shown to be strong push cop-win, we currently lack an explicit example of a graph $\overrightarrow{G}$ with ${\sf c_{sp}}(\overrightarrow{G})>1$. Thus, a compelling direction for future work is to construct such graphs or, in a less likely but intriguing scenario, prove that every graph is cop-win under this model.
@@ 6. Acknowledgments
This research was supported by the IDEX-ISITE initiative CAP 20-25 (ANR-16-IDEX-0001), the International Research Center “Innovation Transportation and Production Systems” of the I-SITE CAP 20-25, and the ANR project GRALMECO (ANR-21-CE48-0004). We thank the anonymous reviewers for their thorough and constructive feedback, which significantly improved both the presentation and the technical clarity of this paper.